OpenStudy (anonymous):

Using the first three terms of the Maclaurin's series, what is the value of cos(Ï€/2)?

OpenStudy (anonymous):

Do you know how to find the MacLaurin/Taylor series of cos(x)?

OpenStudy (anonymous):

No!!

OpenStudy (anonymous):

Here's the general form: MacLaurin series of f(x) centered about x=0: \[f(0) + f'(0)*x + \frac{f''(0)*x^2}{2!} + \frac{f'''(0)*x^3}{3!} + ... + \frac{f^n(0)*x^n}{n!} + ...\] Use this to find the MacLaurin series of cos(x). You should notice a pattern pretty quickly, and since you're only using the first three terms, you want to plug π/2 for x after you get the series.

OpenStudy (anonymous):

Thank you !!

OpenStudy (anonymous):

Sorry, the general term should be f^(n) instead of f^n, because you're taking the n-th derivative of the function at 0. No problem. :)

OpenStudy (anonymous):

O.K.